# ARN: Analogical Reasoning on Narratives

### v3

@@ 6 Conclusions and Future Work
This paper addressed a gap between analogical reasoning theories in cognitive psychology, and analogical reasoning benchmarks in NLP. Our proposed framework studies analogical reasoning over narratives by a three-tier process of extracting narrative elements following narratology, devising corresponding mappings according to analogical reasoning theories, and providing a formal method for quantifying near/far analogical distances to create near/far analogies and disanalogies. The framework provides a solid basis for a binary QA task that enables a systematic study of models’ analogical reasoning abilities across various scenarios. Leveraging data on narratives with high-level proverbial messages, we instantiate the devised task with the Analogical Reasoning on Narratives ($ARN$) benchmark, consisting of 1096 triples of query narratives, analogies, and distractors. Evaluating multiple LLMs on $ARN$ in a zero-shot setting suggests that while models can recognize near analogies, their analogical reasoning performance degrades when detecting far analogies, characterized by the absence of surface mappings. This trend also holds for GPT4.0, performing best on average but dropping to a near-random performance on detecting far analogies in a zero-shot setting. Few-shot prompting with chain-of-thought reasoning enhances models’ performance in far analogies while being detrimental to solving near analogies. Overall, we show that LLMs’ analogical reasoning over narratives lags behind humans, especially on far analogies, which motivates further research on devising computational analogical reasoners on narratives.
@@ 7 Limitations
While our framework showcases its potential within a specific narrative benchmark, extending its application to diverse benchmarks and resources might require additional adaptations. We see additional development of datasets, benchmarks, and discussions around the analogical reasoning capabilities of LLMs extremely useful as they promote generalizability and reusability of models in arbitrary domains models are not trained on. Further, although we used narrative elements to form mappings, alternative approaches like using elements inspired by problem-solving can also be explored (Gick and Holyoak, 1980). Finally, going beyond LLMs, considering other methods inspired by CogPsy and computational models such as neuro-symbolic approaches might offer valuable insights into a more diverse set of models and their analogical reasoning on narratives.
@@ Appendix A Data Collection Details
As we use the ePiC dataset (Ghosh and Srivastava, 2021) as our narrative database, there are certain pre-processings that needed to be done on the original ePiC dataset before they would be ready for the subsequent steps explained in Section 4. We modify all the narratives that do not necessarily reflect the proverb they are supposed to be associated with by rewriting the narratives and keeping the elements of the narratives intact. [^4] This happens since all the narratives are crowdsourced, and limited supervision or structure has been in place when gathering the narratives. We consider two proverbs with the same meaning similar, and hence, the narratives associated with those proverbs are, in turn, analogous. On this ground, we consolidate proverbs and their corresponding narratives that convey the same meaning into a single representative proverb with its associated narratives, which results in the cleaned dataset having 223 unique proverbs and between 10 to 40 narratives associated with each proverb.
@@ C.1 Extracting Narrative Elements
The prompts given to Llama 2, GPT3.5, and GPT4.0 models for the main zero-shot and few-shot experiments consisted of multiple components: 1. the component in which the task is defined and some details are provided about the template that the model should use to return the answers; 2. the component in which the demonstrations are included in the prompt (shown in a lighter color), which is not present in the zero-shot setting and is repeated multiple times in the few-shot setting depending on the specified number of demonstrations. 3. the component that shows the question that the model must solve: narratives can be mapped to each other in terms of the high-level message they strive to convey. These high-level messages can be related to traditions, common knowledge, or moral principles. We call this mapping analogical mapping. Which one of the two narratives (1, 2) can create a better analogical mapping with the query narrative? Answer in the template: {{narrative_x, because narrative_x and query_narrative are …}}
@@ C.3 Macaw, UnifiedQA, and FlanT5 Evaluation
The prompt given to Macaw, UnifiedQA, and FlanT5 models that were utilized in zero-shot experiments was different from the prompt used for GPT3.5, GPT4.0, and Llama 2 models, because of the specific instruction template that was used in their fine-tuning. The content of the prompt is the same as the other prompt; however, to adopt a question-answering setting discussed in Macaw (Tafjord and Clark, 2021), UnifiedQA (Khashabi et al., 2020), and FlanT5 (Chung et al., 2022), we slightly modified the prompt as follows:
@@ D.1 Effect of Scaling Up the Models’ Size on Analogical Reasoning
Figure 7 demonstrates the zero-shot performance of all models on $ARN$ with different numbers of parameters. In line with the prior finding (e.g. Wei et al., 2022), we find that scaling up the models enhances the analogical reasoning capabilities of models. The increase in the performance of models is 10 absolute points, on average, with UnifiedQA gaining the most from scaling up the number of parameters, as high as almost 20 absolute points.
@@ D.2 Distracting Effect of Surface Mappings
Performance of all models on $ARN$ categorized by the type of analogies, disanalogies, and specific surface mappings included in distractors, is demonstrated in Figure 8. We observe that for all the models except for UnifiedQA-11B, including all types of distractors is more distracting than scenarios where some are excluded, especially in far analogies. In other words, each distractor has its own individual effect on the performance and their aggregated effect is stronger than individuals. Also, among different elements, the lower-order mappings that are formed based on the actions of narratives have the most distracting effect on the analogical reasoning abilities of models. However, we see a different trend for UnifiedQA-11B, where including all the distractors does not yield the most distracting effect, and indeed this model has its highest performance when all types of distractors are included.
@@ D.3 Do Models Find Analogies and Similarities Different From Each Other?
While in our main experiments, we explicitly prompt the model to find system mappings and mappings based on proverbial high-level messages, in an additional experiment, we prompted models to look for mere "similarity," with the performance of models in both settings demonstrated in Table 5. We observed that the performance with the explicit instruction to look for analogical mappings is higher compared to the setting without explicit instructions, which is reasonable given the nature of the task and models following the instructions. The relatively small gap between the two settings might be due to the models seeing analogical mappings as more sound compared to surface mappings, which is similar to humans (Gentner et al., 1993). These results suggest that the low performance of the models in Table 2 is not due to the chosen prompt. However, future work can investigate more effective approaches to activate the analogical reasoning abilities of LLMs.

### v4

@@ 6 Conclusions and Future Work
This paper addressed a gap between analogical reasoning theories in cognitive psychology and analogical reasoning benchmarks in NLP. Our proposed framework studies analogical reasoning over narratives by a three-tier process of extracting narrative elements following narratology, devising corresponding mappings according to analogical reasoning theories, and providing a formal method for quantifying near/far analogical distances to create near/far analogies and disanalogies. The framework provides a solid basis for a binary QA task that enables a systematic study of models’ analogical reasoning abilities across various scenarios. Leveraging data on narratives with high-level proverbial messages, we instantiate the devised task with the Analogical Reasoning on Narratives ($ARN$) benchmark, consisting of 1096 triples of query narratives, analogies, and distractors. Evaluating multiple LLMs on $ARN$ in a zero-shot setting suggests that while models can recognize near analogies, their analogical reasoning performance degrades when detecting far analogies, characterized by the absence of surface mappings. This trend also holds for GPT4.0, performing best on average but dropping to a below-random performance on detecting far analogies in a zero-shot setting. Few-shot prompting with Chain-of-Thought reasoning enhances models’ performance in far analogies while being detrimental to solving near analogies. Overall, we show that LLMs’ analogical reasoning over narratives lags behind humans, especially on far analogies, which motivates further research on devising computational analogical reasoners on narratives.
@@ 7 Limitations
While our framework showcases its potential within a specific narrative benchmark, extending its application to diverse benchmarks and resources might require additional adaptations. We see additional development of benchmarks and discussions around the analogical reasoning capabilities of LLMs extremely useful as they promote generalizability and reusability of models in arbitrary domains. Further, although we used narrative elements to form mappings, alternative approaches like using elements inspired by problem-solving can be explored (Gick and Holyoak, 1980). Finally, going beyond LLMs, considering other methods inspired by CogPsy and neuro-symbolic models might offer valuable insights into a more diverse set of models and their analogical reasoning on narratives.
Acknowledgements
This research was supported, in part, by the Army Research Laboratory under contract W911NF-23-2-0183, and by National Science Foundation under Contract No. IIS-2153546. We thank Keith Holyoak and Hongjing Lu for the fruitful discussion around the topic in the initial stages of the study. Moreover, we thank the reviewers and action editors from TACL for their insightful comments and valuable feedback.
@@ Appendix A Data Collection Details
As we use the ePiC dataset (Ghosh and Srivastava, 2022) as our narrative database, there are certain pre-processings that needed to be done on the original ePiC dataset before they would be ready for the subsequent steps explained in Section 4. We modify all the narratives that do not necessarily reflect the proverb they are supposed to be associated with by rewriting the narratives and keeping the elements of the narratives intact. [^4] This happens since all the narratives are crowdsourced, and limited supervision or structure has been in place when gathering the narratives. We consider two proverbs with the same meaning similar, and hence, the narratives associated with those proverbs are, in turn, analogous. On this ground, we consolidate proverbs and their corresponding narratives that convey the same meaning into a single representative proverb with its associated narratives, which results in the cleaned dataset having 223 unique proverbs and between 10 to 40 narratives associated with each proverb.
@@ C.1 Extracting Narrative Elements
The prompts given to Llama 2, GPT3.5, and GPT4.0 models for the main zero-shot and few-shot experiments consisted of multiple components: 1. the component in which the task is defined and some details are provided about the template that the model should use to return the answers; 2. the component in which the demonstrations are included in the prompt (shown in a lighter color), which is not present in the zero-shot setting and is repeated multiple times in the few-shot setting depending on the specified number of demonstrations. 3. the component that shows the question that the model must solve:
narratives can be mapped to each other in terms of the high-level message they strive to convey. These high-level messages can be related to traditions, common knowledge, or moral principles. We call this mapping analogical mapping. Which one of the two narratives (1, 2) can create a better analogical mapping with the query narrative? Answer in the template: {{narrative_x, because narrative_x and query_narrative are …}}
@@ C.3 Macaw, UnifiedQA, and FlanT5 Evaluation
The prompt given to Macaw, UnifiedQA, and FlanT5 models that were utilized in zero-shot experiments was different from the prompt used for GPT3.5, GPT4.0, and Llama 2 models, because of the specific instruction template that was used in their fine-tuning. The content of the prompt is the same as the other prompt; however, to adopt a question-answering setting discussed in Macaw (Tafjord and Clark, 2021), UnifiedQA (Khashabi et al., 2020), and FlanT5 (Chung et al., 2024), we slightly modified the prompt as follows:
@@ D.1 Effect of Scaling Up the Models’ Size on Analogical Reasoning
Figure 7 demonstrates the zero-shot performance of all models on $ARN$ with different numbers of parameters. In line with the prior finding (e.g., Wei et al., 2022a), we found that scaling up the models enhances the analogical reasoning capabilities of models. The increase in the performance of models was 10 absolute points, on average, with UnifiedQA gaining the most from scaling up the number of parameters, as high as almost 20 absolute points.
@@ D.2 Distracting Effect of Surface Mappings
Performance of all models on $ARN$ categorized by the type of analogies, disanalogies, and specific surface mappings included in distractors, is demonstrated in Figure 8. We observed that for all the models except for UnifiedQA-11B, including all types of distractors is more distracting than scenarios where some are excluded, especially in far analogies. In other words, each distractor has its own individual effect on the performance and their aggregated effect is stronger than individuals. Also, among different elements, the lower-order mappings that were formed based on the actions of narratives had the most distracting effect on the analogical reasoning abilities of models. However, we saw a different trend for UnifiedQA-11B, where including all the distractors did not yield the most distracting effect, and indeed the model had its highest performance when all types of distractors were included.
@@ D.3 Do Models Find Analogies and Similarities Different From Each Other?
While in our main experiments, we explicitly prompted the model to find system mappings and mappings based on proverbial high-level messages, in an additional experiment, we prompted models to look for mere "similarity," with the performance of models in both settings demonstrated in Table 5. We observed that the performance with the explicit instruction to look for analogical mappings is higher compared to the setting without explicit instructions, which is reasonable given the nature of the task and models following the instructions. The relatively small gap between the two settings might be due to the models seeing analogical mappings as more sound compared to surface mappings, which is similar to humans (Gentner et al., 1993). These results suggest that the low performance of the models in Table 2 is not due to the chosen prompt. However, future work can investigate more effective approaches to activate the analogical reasoning abilities of LLMs.